# Do Robot Snakes Dream like Electric Sheep? Investigating the Effects of Architectural Inductive Biases on Hallucination

### v2

@@ 4 Experiments and Results
    Closed-book Open-domain QA: NQ-Open (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), TruthfulQA (MC1, MC2, Generative) (Lin et al., 2022), PopQA (Joshi et al., 2017)
@@
[…]
4.1 Investigating Task Biases
At a broader scope that architecture does not appear to influence hallucination, both for task categories and factuality/faithfulness. For example, with like-sized models falling within different categories of their time mixers (self-attention, recurrent layers, hybrid), mean performance on tasks within each grouping is not significantly different nor do they show a particular ordering [^1] Explicit numbers in Table 3, 4 and 5 in Appendix B.. However, with specific tasks, differences start to present themselves, which we describe in our following observations.
O1: Recurrent models struggle with rare knowledge.
Interestingly, recurrent and hybrid models all significantly underwhelm on PopQA, a task that tests for uncommon (thin-tailed) factual knowledge (left plot in Figure 1). For example, FalconMamba achieves an exact-match (EM) score of 0.7 compared to 17.5 by a similar Falcon. Similarly, 2B and 9B RecurrentGemma achieve 7.6 and 16.0 EM respectively, compared to 14.6/21.3 EM for 2B/7B Gemma and 13.6/18.2 by 2B/9B Gemma2, while Jamba (0.4 EM) performs worse than a similarly sized Mixtral-8x7B (31.9 EM).
Previous work (Vardasbi et al., 2023) suggests that hidden states can become dense and difficult to extract information from in the future. Although this is relevant for hallucinations, these concerns are unlikely to apply in this setting given that the prompt consists of only a question and the amount of information to store within the hidden representation is unlikely to lead to it becoming highly dense. Additional works (Dao and Gu, 2024) have instead posited that using recurrent models with additional mechanisms enabling input filtering can enhance performance on information-dense data such as language; however, this is again unlikely to be the case due to the general simplicity of the task. Yet this result is significant as it indicates recurrent models, despite the task’s simplicity, struggle when compared to attention-based equivalents, suggesting a direct link to the token mixing that is utilized. The reason may instead be that recurrent/hybrid models do not learn by explicitly memorizing and acting as a retrieval system (as discussed in Section 3). In this sense, despite potentially observing low-tail information, the learning mechanism may not have led to said information being stored within parameters, making recurrent/hybrid models more likely to produce incorrect information in such settings with higher degrees of uncertainty.
Attention-Only Models
Gemma-2B 49.65 ($\uparrow$ 11.51) 30.25 ($\uparrow$ 13.84) 38.20 ($\downarrow$ -6.92) 38.96 ($\downarrow$ -13.96) 24.93 ($\downarrow$ -4.53) 54.14 ($\uparrow$ 1.60) 35.85 ($\uparrow$ 1.92) 32.56 ($\downarrow$ -4.56)
Gemma-7B 53.81 ($\downarrow$ -5.67) 31.94 ($\uparrow$ 6.79) 27.43 ($\downarrow$ -4.13) 31.35 ($\downarrow$ -5.83) 20.63 ($\uparrow$ 1.04) 43.28 ($\downarrow$ -5.22) 36.48 ($\uparrow$ 2.71) 44.50 ($\downarrow$ -6.04)
Gemma2-2B 58.33 ($\uparrow$ 0.64) 26.73 ($\uparrow$ 10.90) 29.59 ($\uparrow$ 4.83) 32.95 ($\downarrow$ -5.26) 16.04 ($\downarrow$ -0.47) 39.57 ($\uparrow$ 18.96) 35.61 ($\uparrow$ 0.65) 34.96 ($\downarrow$ -0.98)
Gemma2-9B 65.64 ($\uparrow$ 6.31) 23.76 ($\uparrow$ 30.07) 39.36 ($\uparrow$ 3.13) 39.74 ($\downarrow$ -0.70) 21.38 ($\downarrow$ -3.35) 62.06 ($\uparrow$ 6.47) 40.56 ($\uparrow$ 7.55) 46.85 ($\uparrow$ 3.13)
Gemma2-27B 62.03 ($\uparrow$ 12.43) 26.62 ($\uparrow$ 33.58) 47.19 ($\uparrow$ 2.86) 43.04 ($\uparrow$ 5.02) 28.92 ($\downarrow$ -0.74) 68.28 ($\uparrow$ 1.00) 41.54 ($\uparrow$ 13.02) 53.92 ($\uparrow$ 2.36)
LLaMA2-7B 53.63 ($\uparrow$ 4.30) 28.94 ($\uparrow$ 8.99) 37.64 ($\downarrow$ -1.89) 27.58 ($\uparrow$ 3.74) 25.19 ($\uparrow$ 0.69) 51.38 ($\uparrow$ 5.88) 35.14 ($\uparrow$ 4.01) 42.51 ($\downarrow$ -0.12)
LLaMA2-13B 67.80 ($\downarrow$ -3.68) 26.57 ($\uparrow$ 9.72) 39.65 ($\downarrow$ -0.65) 32.17 ($\downarrow$ -3.64) 27.33 ($\downarrow$ -0.23) 62.35 ($\downarrow$ -1.64) 40.83 ($\uparrow$ 0.88) 46.27 ($\downarrow$ -0.49)
LLaMA2-70B 62.34 ($\uparrow$ 10.78) 30.24 ($\uparrow$ 17.11) 47.85 ($\downarrow$ -2.53) 39.48 ($\downarrow$ -8.21) 28.00 ($\downarrow$ -0.55) 66.63 ($\downarrow$ -1.45) 41.74 ($\uparrow$ 5.82) 53.69 ($\downarrow$ -2.34)
LLaMA3-8B 60.76 ($\uparrow$ 10.12) 22.06 ($\uparrow$ 26.76) 41.55 ($\uparrow$ 1.30) 33.52 ($\uparrow$ 3.62) 26.62 ($\downarrow$ -1.60) 60.84 ($\uparrow$ 4.82) 37.60 ($\uparrow$ 10.09) 48.14 ($\uparrow$ 1.60)
LLaMA3-70B 71.78 ($\uparrow$ 8.69) 21.59 ($\uparrow$ 25.60) 48.07 ($\uparrow$ 3.43) 46.38 ($\downarrow$ -7.73) 28.91 ($\downarrow$ -1.68) 69.57 ($\uparrow$ 0.99) 45.92 ($\uparrow$ 6.23) 54.80 ($\uparrow$ 2.81)
Mistral-7B 60.48 ($\uparrow$ 3.36) 28.39 ($\uparrow$ 16.85) 41.24 ($\uparrow$ 5.21) 32.03 ($\uparrow$ 1.92) 26.77 ($\downarrow$ -0.77) 58.59 ($\uparrow$ 6.78) 38.47 ($\uparrow$ 4.67) 47.42 ($\uparrow$ 4.17)
Mixtral-8x7B 73.51 ($\downarrow$ -0.15) 26.84 ($\uparrow$ 17.60) 48.99 ($\uparrow$ 4.33) 40.66 ($\downarrow$ -3.90) 27.91 ($\downarrow$ -1.07) 68.35 ($\uparrow$ 0.49) 46.16 ($\uparrow$ 3.89) 55.15 ($\uparrow$ 3.50)
Recurrent and Hybrid Models
RecurrentGemma-2B 52.88 ($\uparrow$ 2.33) 33.43 ($\uparrow$ 3.12) 27.36 ($\downarrow$ -1.85) 30.66 ($\downarrow$ -1.77) 18.41 ($\uparrow$ 2.86) 42.97 ($\uparrow$ 3.63) 36.55 ($\downarrow$ -0.14) 31.46 ($\uparrow$ 1.39)
RecurrentGemma-9B 55.75 ($\downarrow$ -1.67) 31.67 ($\uparrow$ 12.96) 36.79 ($\uparrow$ 2.03) 36.57 ($\downarrow$ -6.66) 22.99 ($\uparrow$ 1.84) 51.25 ($\uparrow$ 13.36) 37.61 ($\downarrow$ -0.17) 42.96 ($\uparrow$ 3.28)
FalconMamba-7B 55.80 ($\uparrow$ 0.73) 42.97 ($\downarrow$ -1.19) 39.94 ($\downarrow$ -0.43) 23.76 ($\downarrow$ -0.85) 23.89 ($\downarrow$ -0.09) 61.85 ($\downarrow$ -1.80) 35.92 ($\downarrow$ -0.26) 47.02 ($\downarrow$ -0.38)
O2: Recurrent models can better follow contexts by relying less on memory.
There are also cases where recurrent models consistently outperform attention equivalents, namely on MemoTrap (right plot in Figure 1). Here, the LLM is prompted to complete a well-known proverb with an ending that deviates from the commonly used ending, testing for over-reliance on knowledge memorized from a training corpus. This suggests that although recurrent models might struggle with long-tail knowledge, they are less prone to ignoring contextual cues not stored within parametric memory, or they more model the dynamics based on the context as opposed to acting as a retrieval system. Noticeable drops exist from FalconMamba (58.8) to Falcon (38.9) and from RecurrentGemma 2B/9B to Gemma2 2B/9B (53.7/43.4 vs 42.7/34.4), suggesting that attention can lead to more hallucinations. It further reveals a benefit of recurrent layers; by not memorizing information directly within parameters, there may be a reduced tendency to repeat previous observations (Jelassi et al., 2024) and greater focus on the context.
O3: Scale is required for emergent qualities.
However, some scale to the data and model remain important. Note that Mamba, Finch and Pythia fail to draw the same distinctions as the other models despite sharing the same vocabulary and training data. However, given the scale of training for these (300B tokens compared to >1T tokens for other models) and the fact that the discussed tasks relate directly to information memorization, it is possible that these phenomena fail to emerge at this data scale. This is particularly evidenced by how these models perform better with size on MemoTrap, which is designed in a way such that larger models should normally perform worse.
4.2 On the role of Instruction-Tuning.
However, changes in faithfulness are generally positive for pure attention models (shown in teal) and negative for models with recurrent layers (shown in magenta), whereas changes in factuality show no consistent trends. This suggests that instruction-tuning for LLMs, as initially proposed with self-attention based models, enables particular properties which are no longer relevant after the addition of recurrent layers. Their inductive biases therefore can have a particular effect on specific forms of hallucinations emerging, with instruction-tuning being a manner to mitigate these.
Recalling the links between attention and retrieval from Section 3, the process of instruction-tuning can be understood as shifting the projection weights towards a new space through the additional fine-tuning (which is of itself simply an additional step of language modeling). This leads to the over-writing of some of the pre-training corpus with that used for fine-tuning, effectively replacing the keys and values from the underlying retrieval storage. While effects are ambiguous on factuality, as some potentially relevant facts can be removed, faithfulness often improves from this process, as the fine-tuning corpus is more likely to contain information relevant to context following and thereby make it easier for the attention mechanism to retrieve relevant information for such a purpose. In addition to highlighting how enhancing the ability to follow instructions does not necessarily lead to factually correct results, another conclusion stemming from the fact that recurrent models generally observe no benefits in faithfulness after instruction-tuning is that such models’ inductive biases render them inherently more faithful or the process of making them more faithful is distinct from self-attention.
4.3 Impact of Model Size on Hallucinations.
Immediately, we observe that increased size leads to improvements in factuality regardless of model. This is partially expected, considering the range of work that has suggested links between size and the amount of factual knowledge that can be encoded within parametric space. However, faithfulness is different. Recurrent and hybrid architectures tend to saturate in terms of faithfulness at a small size whereas pure attention models do not. For example, Mamba models stagnate in faithfulness at around 130M, whereas Pythia continues to increase in this score up to billions of parameters. Finch and RecurrentGemma also show this while attention-based models do not. Another observation is the fact that even the smallest Mamba are nearly as faithful as LLaMA2-7B and Pythia-1.4B and despite models eventually catching up in terms of faithfulness, having such small models perform on par indicates a meaningful manner in which their hallucinations may differ. To further illustrate this, we provide an additional depiction of a performance to scale metric, defined as
[…]
5 Discussion
Hallucination Mitigation for Different Architectures.
Hallucinations motivate the need to both detect and mitigate them, particularly when they can escalate to have adverse effects. Existing techniques may be insufficient, as they contain potential implicit biases that can fundamentally conflict with the manners in which hallucinations may differ between models. Facts such as sequence models saturating in faithfulness in a manner not resolved through instruction-tuning indicate that existing mitigation techniques can differ in their effectiveness due to a bias towards specific architectures.
This means that current techniques also further deserve a look, in particular in terms of their effectiveness across the different inductive biases of models. While some methods evidently might be specifically catered to specific models, e.g. using attention (Li et al., 2023b; Zhang et al., 2024; Chuang et al., 2024), other methods may implicitly also possess biases due to the underlying assumptions being made or specific requirements that may be unevenly addressed through different architectures, such as the reliance on data-refinement or additional context (Shi et al., 2024).
Inherent Faithfulness and Factuality.
Nevertheless, these results hint that architectures can provide an avenue through which desiderata can be attained. SSMs present an option which in particular appears suited for settings which require utilizing information provided within a context, given their use within control theory to model system dynamics. With language, this can be analogous to using the fixed-size representation of the previous context for control, which may mean that such models can remain more faithful as the context remains the source that guides the generation. In comparison, attention-based models have been shown to act as retrieval systems that retrieve tokens based on high similarity with memorized information; hence, when a large difference exists between the context and what has been stored from training, breakdowns in faithfulness can occur.
6 Conclusion
7 Limitations
Evaluating faithfulness and factuality.
Limits on tasks.
8 Ethical Concerns
9 Acknowledgements
Jerry Huang received partial support in the form of a National Science and Engineering Research Council (NSERC) Canada Graduate Scholarship, a Fonds de Recherche du Québec Nature et Technologies (FRQNT) Training Scholarship, and a Hydro-Québec Excellence Scholarship during the course of this work. Sarath Chandar is supported by a Canada CIFAR AI Chair, the Canada Research Chair in Lifelong Machine Learning, and a NSERC Discovery Grant.
Appendix A Technical Implementation Details
A.1 Models and Datasets
A.2 Experimental Setup
A.3 Datasets
  1. 1)
    Closed-book Open-domain QA: NQ-Open (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), TruthfulQA (MC1, MC2, Generative) (Lin et al., 2022), PopQA (Joshi et al., 2017)
  2. 2)
    Summarization: XSum (Narayan et al., 2018), CNN/DM (See et al., 2017)
  3. 3)
  4. 4)
  5. 5)
  6. 6)
    Fact Checking: FEVER (Thorne et al., 2018)

### v5

@@ 4 Experiments and Results
    Closed-book Open-domain QA: NQ-Open (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), TruthfulQA (Lin et al., 2022), PopQA (Mallen et al., 2023)
  2. 2)
    Summarization: XSum (Narayan et al., 2018), CNN/DM (See et al., 2017)
  3. 3)
  4. 4)
  5. 5)
    Hallucination Detection: FaithDial (Dziri et al., 2022a), HaluEval (Li et al., 2023a), True-False (Azaria and Mitchell, 2023)
  6. 6)
    Fact Checking: FEVER (Thorne et al., 2018)
[…]
4.1 Investigating Task Biases
Our first interest is to verify whether specific choices in architecture can lead to highly evident patterns in performance [^1] Numbers provided in Table 4, 5 and 8 in Appendix B.. While direct differences can be difficult to quantify due to differences in how each model is trained, some pattern are consistent, described as follows.
O1: Sequence models miss rare knowledge.
Interestingly, recurrent and hybrid models all significantly underwhelm on PopQA, a task that tests for uncommon knowledge (left plot in Figure 1). For example, FalconMamba achieves an exact-match (EM) score of 0.7 compared to 17.5 by a similar Falcon. Similarly, 2B and 9B RecurrentGemma achieve 7.6 and 16.0 EM compared to 14.6/21.3 EM for 2B/7B Gemma and 13.6/18.2 by 2B/9B Gemma2, while Jamba (0.4 EM) performs worse than a similarly sized Mixtral (31.9 EM).
Previous work (Vardasbi et al., 2023) suggests that hidden states can become dense and difficult to extract information from. However, the brevity and simplicity of the prompt here begs the question whether the hidden representation has the opportunity to become significantly dense. Additional works (Dao and Gu, 2024) instead posit that using recurrent models with additional mechanisms enabling input filtering can enhance performance on domains such as language; however, the relevance of this claim here is again questionable. Yet this result is significant as it indicates recurrent models, despite the task’s simplicity, struggle when compared to attention-based equivalents, suggesting a link with the sequence mixing. Instead, the reason may be that recurrent/hybrid models do not learn by explicitly memorizing and acting as a retrieval system (as discussed in Section 3), where the learning mechanism may not have led to said information being stored within parameters.
Attention-Only Models
Gemma-2B 49.65 ($\uparrow$ 11.51) 30.25 ($\uparrow$ 13.84) 38.20 ($\downarrow$ -6.92) 38.96 ($\downarrow$ -13.96) 24.93 ($\downarrow$ -4.53) 54.14 ($\uparrow$ 01.60) 35.85 ($\uparrow$ 01.92) 32.56 ($\downarrow$ -4.56)
Gemma-7B 53.81 ($\downarrow$ -5.67) 31.94 ($\uparrow$ 06.79) 27.43 ($\downarrow$ -4.13) 31.35 ($\downarrow$ -5.83) 20.63 ($\uparrow$ 01.04) 43.28 ($\downarrow$ -5.22) 36.48 ($\uparrow$ 02.71) 44.50 ($\downarrow$ -6.04)
Gemma2-2B 58.33 ($\uparrow$ 00.64) 26.73 ($\uparrow$ 10.90) 29.59 ($\uparrow$ 04.83) 32.95 ($\downarrow$ -5.26) 16.04 ($\downarrow$ -0.47) 39.57 ($\uparrow$ 18.96) 35.61 ($\uparrow$ 00.65) 34.96 ($\downarrow$ -0.98)
Gemma2-9B 65.64 ($\uparrow$ 06.31) 23.76 ($\uparrow$ 30.07) 39.36 ($\uparrow$ 03.13) 39.74 ($\downarrow$ -0.70) 21.38 ($\downarrow$ -3.35) 62.06 ($\uparrow$ 06.47) 40.56 ($\uparrow$ 07.55) 46.85 ($\uparrow$ 03.13)
Gemma2-27B 62.03 ($\uparrow$ 12.43) 26.62 ($\uparrow$ 33.58) 47.19 ($\uparrow$ 02.86) 43.04 ($\uparrow$ 05.02) 28.92 ($\downarrow$ -0.74) 68.28 ($\uparrow$ 01.00) 41.54 ($\uparrow$ 13.02) 53.92 ($\uparrow$ 02.36)
LLaMA2-7B 53.63 ($\uparrow$ 04.30) 28.94 ($\uparrow$ 08.99) 37.64 ($\downarrow$ -1.89) 27.58 ($\uparrow$ 03.74) 25.19 ($\uparrow$ 00.69) 51.38 ($\uparrow$ 05.88) 35.14 ($\uparrow$ 04.01) 42.51 ($\downarrow$ -0.12)
LLaMA2-13B 67.80 ($\downarrow$ -3.68) 26.57 ($\uparrow$ 09.72) 39.65 ($\downarrow$ -0.65) 32.17 ($\downarrow$ -3.64) 27.33 ($\downarrow$ -0.23) 62.35 ($\downarrow$ -1.64) 40.83 ($\uparrow$ 00.88) 46.27 ($\downarrow$ -0.49)
LLaMA2-70B 62.34 ($\uparrow$ 10.78) 30.24 ($\uparrow$ 17.11) 47.85 ($\downarrow$ -2.53) 39.48 ($\downarrow$ -8.21) 28.00 ($\downarrow$ -0.55) 66.63 ($\downarrow$ -1.45) 41.74 ($\uparrow$ 05.82) 53.69 ($\downarrow$ -2.34)
LLaMA3-8B 60.76 ($\uparrow$ 10.12) 22.06 ($\uparrow$ 26.76) 41.55 ($\uparrow$ 01.30) 33.52 ($\uparrow$ 03.62) 26.62 ($\downarrow$ -1.60) 60.84 ($\uparrow$ 04.82) 37.60 ($\uparrow$ 10.09) 48.14 ($\uparrow$ 01.60)
LLaMA3-70B 71.78 ($\uparrow$ 08.69) 21.59 ($\uparrow$ 25.60) 48.07 ($\uparrow$ 03.43) 46.38 ($\downarrow$ -7.73) 28.91 ($\downarrow$ -1.68) 69.57 ($\uparrow$ 00.99) 45.92 ($\uparrow$ 06.23) 54.80 ($\uparrow$ 02.81)
Mistral-7B 60.48 ($\uparrow$ 03.36) 28.39 ($\uparrow$ 16.85) 41.24 ($\uparrow$ 05.21) 32.03 ($\uparrow$ 01.92) 26.77 ($\downarrow$ -0.77) 58.59 ($\uparrow$ 06.78) 38.47 ($\uparrow$ 04.67) 47.42 ($\uparrow$ 04.17)
Mixtral-8x7B 73.51 ($\downarrow$ -0.15) 26.84 ($\uparrow$ 17.60) 48.99 ($\uparrow$ 04.33) 40.66 ($\downarrow$ -3.90) 27.91 ($\downarrow$ -1.07) 68.35 ($\uparrow$ 00.49) 46.16 ($\uparrow$ 03.89) 55.15 ($\uparrow$ 03.50)
Recurrent and Hybrid Models
RecurrentGemma-2B 52.88 ($\uparrow$ 02.33) 33.43 ($\uparrow$ 03.12) 27.36 ($\downarrow$ -1.85) 30.66 ($\downarrow$ -1.77) 18.41 ($\uparrow$ 02.86) 42.97 ($\uparrow$ 03.63) 36.55 ($\downarrow$ -0.14) 31.46 ($\uparrow$ 01.39)
RecurrentGemma-9B 55.75 ($\downarrow$ -1.67) 31.67 ($\uparrow$ 12.96) 36.79 ($\uparrow$ 02.03) 36.57 ($\downarrow$ -6.66) 22.99 ($\uparrow$ 01.84) 51.25 ($\uparrow$ 13.36) 37.61 ($\downarrow$ -0.17) 42.96 ($\uparrow$ 03.28)
FalconMamba-7B 55.80 ($\uparrow$ 00.73) 42.97 ($\downarrow$ -1.19) 39.94 ($\downarrow$ -0.43) 23.76 ($\downarrow$ -0.85) 23.89 ($\downarrow$ 0-0.09) 61.85 ($\downarrow$ -1.80) 35.92 ($\downarrow$ -0.26) 47.02 ($\downarrow$ -0.38)
O2: Recurrent models may rely less on memory, leading to more reliance on context.
There are also cases where recurrent models consistently outperform attention equivalents, namely on MemoTrap (right plot in Figure 1). Here, the LLM is prompted to complete a well-known proverb with an ending that deviates from the commonly used ending, testing for over-reliance on knowledge memorized from a training corpus. This suggests that although recurrent models might struggle with long-tail knowledge, they are less prone to ignoring contextual cues not stored within parametric memory, or they more model the dynamics based on the context as opposed to acting as a retrieval system. Noticeable drops exist from FalconMamba (58.8) to Falcon (38.9) and from RecurrentGemma 2B/9B to Gemma2 2B/9B (53.7/43.4 vs 42.7/34.4), suggesting that attention can lead to more hallucinations. It further reveals a benefit of recurrent layers; by not memorizing information directly, there may be a reduced tendency to repeat previous observations (Jelassi et al., 2024) and greater focus on the context.
O3: Scale is required for emergent qualities.
However, some scale to the data and model remain important. Mamba, Finch and Pythia fail to draw the same distinctions as the other models, but given the scale of training for these (300B tokens compared to >1T tokens for other models) and the fact that the discussed tasks relate directly to information memorization, it is possible that these phenomena fail to emerge at this data scale. This is particularly evidenced by how these models all perform better with size on MemoTrap, which is designed in a way such that larger models should normally perform worse.
4.2 On the role of Instruction-Tuning.
However, changes in faithfulness are generally positive for pure attention models and negative for models with recurrent layers, whereas changes in factuality show no consistent trends. This suggests that instruction-tuning for LLMs may enable particular learning patterns which become irrelevant when using recurrent layers. Their inductive biases therefore can have a particular effect on specific forms of hallucinations emerging, with instruction-tuning being a manner to mitigate these.
Recalling from Section 3, instruction-tuning can be understood as shifting the projection weights towards a new space through the additional fine-tuning (which is of itself simply an additional step of language modeling). This leads to the over-writing of some of the pre-training corpus with that used for fine-tuning, effectively replacing the keys and values from the underlying retrieval storage. While effects are ambiguous on factuality, as some potentially relevant facts can be removed, faithfulness often improves from this process, as the fine-tuning corpus is more likely to contain information relevant to context following and thereby make it easier for the attention mechanism to retrieve relevant information for such a purpose. In addition to highlighting how enhancing the ability to follow instructions does not necessarily lead to factually correct results, another conclusion stemming from the fact that recurrent models generally observe no benefits in faithfulness after instruction-tuning is that such models’ inductive biases render them inherently more faithful or the process of making them more faithful is distinct from self-attention.
4.3 Impact of Model Size on Hallucinations.
Immediately, we observe that increased size leads to improvements in factuality regardless of model. This is partially expected, considering the range of work that suggest a link between size and the amount of factual knowledge that can be encoded within parametric space. However, faithfulness is different. Recurrent and hybrid architectures tend to saturate at a small size whereas pure attention models do not. For example, Mamba stagnates in faithfulness at around 130M parameter, whereas Pythia continues to increase up to the billions. Finch and RecurrentGemma also exhibit similar behavior. Furthermore, even the smallest Mamba are equally faithful as a 7B LLaMA2 and 1.4B Pythia, indicating a meaningful manner in which their hallucinations may differ. To further illustrate, we show a performance to scale metric, defined as
[…]
Category Baseline (LLaMA3-8B) 25% Reset 50% Reset 25% Replaced 50% Replaced
Hallucination Detection 60.76 ($\uparrow$ 10.12) 50.27 ($\uparrow$ 03.24) 44.25 ($\uparrow$ 08.53) 53.64 ($\uparrow$ 00.91) 53.29 ($\uparrow$ 0.33)
Instruction Following 22.06 ($\uparrow$ 26.76) 14.29 ($\uparrow$ 6.85) 13.11 ($\uparrow$ 6.50) 20.81 ($\downarrow$ -0.92) 20.18 ($\uparrow$ 0.32)
Closed-Book QA 41.55 ($\uparrow$ 01.30) 36.79 ($\downarrow$ -0.09) 31.69 ($\downarrow$ -0.07) 36.26 ($\downarrow$ -0.42) 33.74 ($\uparrow$ 0.14)
Reading Comprehension 33.52 ($\uparrow$ 03.62) 30.87 ($\uparrow$ 01.85) 29.65 ($\uparrow$ 02.25) 32.16 ($\uparrow$ 00.51) 30.68 ($\uparrow$ 0.03)
Summarization 26.62 ($\downarrow$ -1.60) 21.12 ($\uparrow$ 01.06) 20.86 ($\uparrow$ 01.67) 23.84 ($\uparrow$ 00.54) 22.59 ($\uparrow$ 0.79)
Fact-Checking 60.84 ($\uparrow$ 04.82) 54.93 ($\uparrow$ 02.69) 50.82 ($\uparrow$ 03.17) 55.45 ($\downarrow$ -1.06) 49.66 ($\uparrow$ 2.16)
Faithfulness 37.60 ($\uparrow$ 10.09) 31.40 ($\uparrow$ 04.29) 30.30 ($\uparrow$ 04.48) 35.74 ($\uparrow$ 00.17) 34.32($\uparrow$ 0.13)
Factuality 48.14 ($\uparrow$ 01.60) 43.16 ($\uparrow$ 00.83) 38.98 ($\uparrow$ 01.60) 42.85 ($\downarrow$ -0.31) 39.08($\uparrow$ 0.56)
Table 2: Controlling for the pre-training and post-training data as well as the presence of a channel mixer. Baseline is a LLaMA3-8B model. Colored values in brackets represent change in category performance after an additional phase of supervised instruction tuning. Note that the SFT dataset for the baseline differs from other models.
4.4 Controlling for Data/Model Differences
Models are often trained on varying amounts of data, both for pre-training and fine-tuning, while also potentially possessing additional architectural differences in components, such as a channel mixer that exchanges information within each individual token. These can each influence the representation and potentially have a downstream effect on hallucination. To verify whether or not this may be the case, we control for each of these differences in order to quantify the direct effects that exist from replacing attention with sequence layers.
We inspire ourselves from Wang et al. (2024); we start with a pre-trained base model and replace some attention with sequence layers, namely Mamba, while retaining a similar parameter count and keeping the channel mixing Gated MLPs. For fair comparison, we also consider a scenario where the replaced layers are simply re-initialized. These layers are then trained using a standard next-token prediction objective. Specifically, we use a 8B LLaMA3 model and replace/re-initialize 25% or 50% of the attention layers in the model. These layers are then trained, providing us with a synthetically pre-trained model for which both the pre-training data and the channel mixer are controlled for, enabling us to directly compare the effects of the use of sequence layers compared to attention layers. As a next step, we also perform an additional end-to-end instruction tuning. Thus we control for:
  1. (1)
    The pre-training phase, through the base model and dataset used for re-training.
  2. (2)
    The instruction fine-tuning process, through the use of the same instruction dataset mix.
  3. (3)
    Additional architectural components through when they are frozen and trained.
We use a filtered SlimPajama (Soboleva et al., 2023) to re-train layers with a sequence length of 4096. For the instruction tuning phase, we use the same datasets as Wang et al. (2024) (Chen et al., 2024a; BAAI, 2024; Teknium, 2023). We follow hyper-parameters provided by Wang et al. (2024).
Results in Table 2 demonstrate that our previous observations persist. Replacing attention layers with sequence layers and then re-training appears to regain the losses from the initial baseline model, whereas simply resetting the attention layers still appears to observe a noticeable gap. Findings regarding the effects of instruction fine-tuning hold as well, which is that pure-attention models observe a much greater increase in faithfulness relative to factuality, while hybrid models do not.
5 Discussion
Mitigating Hallucinations in LLMs.
Facts such as sequence models saturating in faithfulness in a manner not resolved through instruction-tuning indicates that existing techniques for hallucination mitigation can differ in their effectiveness due to a bias towards specific architectures. Current techniques therefore deserve a look, in particular in terms of their effectiveness across the different inductive biases of models. While some methods evidently might be specifically catered to specific models, e.g. using attention (Li et al., 2023b; Zhang et al., 2024; Chuang et al., 2024), other methods may implicitly also possess biases due to the underlying assumptions being made or specific requirements that may be unevenly addressed through different architectures, such as the reliance on data-refinement or additional context (Shi et al., 2024).
Inherent Faithfulness and Factuality.
6 Conclusion
7 Limitations
Evaluating faithfulness and factuality.
Limits on tasks.
8 Ethical Concerns
9 Acknowledgements
Jerry Huang was supported by a National Science and Engineering Research Council (NSERC) Canada Graduate Scholarship, a Fonds de Recherche du Québec Nature et technologies (FRQNT) Training Scholarship and a Hydro-Québec Excellence Scholarship. Sarath Chandar is supported by a Canada CIFAR AI Chair, the Canada Research Chair in Lifelong Machine Learning and a NSERC Discovery Grant.
Appendix A Technical Implementation Details
A.1 Models and Datasets
A.2 Experimental Setup
A.3 Datasets
  1. 1)
    Closed-book Open-domain QA: NQ-Open (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), TruthfulQA (MC1, MC2, Generative) (Lin et al., 2022), PopQA (Mallen et al., 2023)
@@ A.3 Datasets
For tasks, a higher score (ranging from 0 to 100) indicates better performance. These are
  • EM: HaluEval (all sets), PopQA, TruthfulQA (MC1 and MC2), SQuADv2, NaturalQuestions (NQ-Open), TriviaQA
  • Accuracy: MemoTrap, TruthfulQA (Gen), RACE, IfEval
  • Rouge-L: CNNDM, XSum, FaithDial, TrueFalse, Fever10